OpenStudy (anonymous):

Suppose the diameter of a circle is 10. What is its area?

OpenStudy (kendricklamar2014):

Area of a Circle:\[A = \pi r^2\]

OpenStudy (kendricklamar2014):

To find the Radius:\[Diameter = 10/2 = 5\]

OpenStudy (kendricklamar2014):

So now we plug in what we know:\[A=\pi5^2\]

OpenStudy (kendricklamar2014):

\[Area = 78.54\]

OpenStudy (anonymous):

it was 25 pi

OpenStudy (kendricklamar2014):

\[5^2 = 25\]

OpenStudy (kendricklamar2014):

next time @flowers4547 post the answer choices as well

OpenStudy (anonymous):

there were no choice options if there was I would say them I'm not that stupid! and I'm worried about my state tests tomorrow.